### v2

# An Overview of Cyber Security Funding for Open Source Software

Abstract
Context: Many open source software (OSS) projects need more human resources for maintenance, improvements, and sometimes even their survival. This need allegedly applies even to vital OSS projects that can be seen as being a part of the world’s critical infrastructures. To address this resourcing problem, new funding instruments for OSS projects have been established in recent years. Objectives: The paper examines two such funding bodies for OSS and the projects they have funded. The focus of both funding bodies is on software security and cyber security in general. Methods: The methodology is based on qualitative thematic analysis. Results: Particularly OSS supply chains, network and cryptography libraries, programming languages, and operating systems and their low-level components have been funded and thus seen as critical in terms of cyber security by the two funding bodies. Conclusions: In addition to the qualitative results presented, the paper makes a contribution by connecting the research branches of critical infrastructure and sustainability of OSS projects. A further contribution is made by connecting the topic examined to recent cyber security regulations. Finally, an important argument is raised that neither cyber security nor sustainability alone can entirely explain the rationales behind the funding decisions made by the two bodies.
keywords:
criticality, critical infrastructure, software security, funding, grants, tragedy of the commons, regulations
1 Introduction
The paper presents a qualitative analysis of open source software projects funded through the Sovereign Tech Agency (STA) and a fund established by the Open Source Security Foundation (OpenSSF). The latter is tied to the larger Linux Foundation, while the former is operated and financed by the Germany’s central government.
For the paper’s purposes, the STA is interesting because it explicitly acknowledges that OSS is an important part of the world’s critical infrastructures. In other words, the Agency “invests in open digital infrastructure”, which refers to “foundational technologies that enable the creation of other software”, and regarding criticality, people often do not “notice how much our lives depend on digital infrastructure until it stops working(STA, 2024i, ). The OpenSSF’s funding instrument is based on a similar rationale; it seeks to protect “society by catalyzing sustainable security improvements to the most critical open source software projects” (OpenSSF, 2024a, ). Criticality is again one of the main keywords.
To the best of the authors’ knowledge, the paper is the first to examine these two funding bodies and the OSS projects they have funded from a cyber security perspective. In addition to the novelty of the paper’s topic, the paper contributes to two research branches. The first branch is the research on critical infrastructure. As will be argued in the opening Section 2, this branch has often overlooked software’s role in critical infrastructures. In particular, there appears to be very limited existing work on the relationship between OSS and critical infrastructures.
The second research branch is about the survivability or sustainability of OSS projects (Alami et al.,, 2024; Avelino et al.,, 2019; Chengalur-Smith et al.,, 2010; Decan et al.,, 2020; Tulili et al.,, 2025). This branch has focused on a question of what makes open source software projects survive over time. An example of a typical explanatory factor considered in the literature is a project’s ability to attract new developers. However, regarding software quality, including software security with it, human resources are not necessarily what matter the most; software’s growing size, complexity, and technical debt may matter more (Alami et al.,, 2024). Also financial resources may well matter for the security of open source software.
To this and other ends, recent research has also started to examine funding of OSS projects. The paper contributes to this recent OSS funding research and its calls for further research and better engagement (Osborne,, 2024; Kula and Treude,, 2025; Sharma,, 2023; Tsakpinis and Pretschner,, 2025). In this regard, it should be emphasized that the STA and OpenSSF are by no means the only funding bodies available for OSS projects today. In addition to so-called micro-donations and crowd-sourcing endeavors (Osborne et al.,, 2024; Tsakpinis and Pretschner,, 2025), there are new funding initiatives for companies to support open source software projects (FCF,, 2025; OSTIF,, 2025), philanthropic programs (CZI,, 2025; NLnet Foundation,, 2025), and funding bodies that align with civil liberties (OTF,, 2025). Given the overall funding landscape, including with respect to scientific research (Braun,, 1998; Cocos,, 2020), these and other funding schemes for OSS projects are strategic in their nature; they focus on some particular problem areas. With respect to the STA and OpenSSF, the strategic problem area is cyber security in general and software security in particular. By implication, the paper’s particular focus on the STA and OpenSSF also connects the survivability research branch with the critical infrastructure branch.
Money and cyber security have an intrinsic relationship. In fact, an argument has often been raised that among the foremost reasons for the world’s cyber insecurity is a lack of incentives for companies, societies, and even individuals, including software engineers (Anderson and Moore,, 2006; Halderman,, 2010). This lack of incentives manifests itself in observations that security is not prioritized in the software industry because it is seen as a cost and not a feature (Ryan et al.,, 2023). Although techniques have been proposed to incentivize and prioritize software security (Weir et al.,, 2023), economic factors, such as cost-effectiveness and return-on-investment, have continued to hamper the adoption of new security solutions for software engineering (Ramaj et al.,, 2024). These obstacles are pronounced in the OSS domain already due to the so-called tragedy of the (digital) commons (Greco and Floridi,, 2004; OpenSSF, 2024e, ). Although this classical tragedy is a contested concept (Frischmann et al.,, 2019), it serves to illustrate a common reasoning; because OSS is openly and publicly available, there is even a less incentive to financially invest in OSS and its security. This tragedy has been seen as a primary reason behind the new funding initiatives (Sharma,, 2023). The tragedy of the (digital) commons further demonstrates that the fundamental issue is not only about the lack of incentives but actually about misplaced incentives. Because some OSS components are nowadays used in practically all commercial software projects, including those addressing the world’s critical infrastructures, the misplaced incentive to not invest in the security of OSS manifests itself in the world’s overall cyber insecurity, including with respect to digital infrastructures that are vital to the functioning of societies.
The paper’s remainder is structured into five sections. The opening Section 2 motivates the background in more detail with respect to the concept of critical infrastructure and its relation to software. Given this elaboration, the subsequent Section 3 presents and motivates the research questions. Then, Section 4 elaborates the empirical material and the qualitative analysis techniques used to examine the material. Results are presented in Section 5. A concluding discussion is presented in the final Section 6. With these remaining five sections, the paper paves the way for further research on an emerging, relevant new research area, providing also a few insights relevant to practitioners, including OSS projects seeking funding, the two funding bodies’ staff, and policy-makers more generally.
2 Motivation
There is a long-standing debate regarding the concept of infrastructure and what is recognized as critical (Eckhardt and Kotovskaia,, 2023; Katina and Hester,, 2013; Zebrowski and Sage,, 2017). The jurisprudence of the European Union (EU) provides a good way to briefly illustrate typical definitions and some associated problems. Thus, the EU’s now deprecated Directive 2008/114/EC defined critical infrastructure as “an asset, system or part thereof” that is located in any given member state and “which is essential for the maintenance of vital societal functions, health, safety, security, economic or social well-being of people, and the disruption or destruction of which would have a significant impact”. Although this definition aligns with the classical concept of an asset, which is essential in traditional information security, the societal focus enlarges the scope from mere information to broader assets with vital society-wide functions.
In this regard, the definition aligns with theoretical conceptualizations to distinguish cyber security from information security (von Solms and van Niekerk,, 2013). The definition is also essentially sectoral (Lehto,, 2022). However, the deprecated directive only considered the energy and transport sectors as critical infrastructure. To address this limitation and other problems, the new Directive (EU) 2022/2557 uses a similar definition but enumerates eleven sectors as potentially critical. Although a selection of critical sectors is left to the member states, the eleven potential sectors highlight further conceptual expansion; among the potentially critical sectors are energy, transport, banking and finance, healthcare, water and food supplies, space technologies, and publication administration. In terms of information technology, many of the Internet’s core technologies are also listed as potentially critical, including domain name system (DNS) providers, top-level domain name registries, cloud computing services and data centers, and content delivery networks. The eleven critical sectors are identical to those specified in Directive (EU) 2022/2555, which is known as the NIS2 directive.
These EU laws have received some criticism in that too many sectors might be covered, expertise might be lacking in some sectors, interoperability and compliance burden may become issues, fragmentation might occur because of sectoral differences and the national adoptions by the member states, and sectors defined as non-critical might receive less resources and effort (Besiekierska,, 2023; Ruohonen,, 2024; Vandezande,, 2024; Veigurs et al.,, 2024). A related critical point is that software has received only a limited explicit attention in the laws. Although the EU’s jurisprudence recognizes information technology as a distinct critical infrastructure of its own, it is important to emphasize that also all of the other critical infrastructures noted are dependent on software. Neither banking nor a central government can nowadays function without software. Nor are there space technologies without software. In fact, it can be argued that a new “software-defined” paradigm is emerging—or has already emerged, and this paradigm also amplifies security risks due to software’s insecurity and other problems (Wu,, 2024). This point justifies also the STA’s and OpenSSF’s overall rationales.
The literature on critical infrastructure has traditionally touched software through its focus on industrial control systems, supervisory control and data acquisition (SCADA) systems, and related technical solutions (Lehto,, 2022). However, this framing is arguably too narrow because also numerous non-specialized software components are underneath the world’s critical infrastructures.
Software is also a good example on how different critical infrastructures are often interlinked to each other through dependencies (Katina and Hester,, 2013; Rinaldi et al.,, 2001). In other words, two or more critical infrastructures may be dependent on each other through a shared software component used in each infrastructure. Then, a single vulnerability in the shared software component may impact several critical infrastructures. Such vulnerabilities have supposedly started to increasingly affect also industrial control and SCADA systems because vendors of such systems have moved toward using commercial off-the-shelf (COTS) operating systems, software architectures, and network protocols (Alanazi et al.,, 2023; Igure et al.,, 2006). Although definite answers are difficult to give, many such COTS solutions are supposedly using also OSS components. The same point applies to the Internet’s critical infrastructures; whether is operating systems, DNS, border gateway protocol (BGP), firewalls, email, or web servers and web browsers, open source software is either fully or partially used to power the infrastructures.
3 Research Questions
The discussion in the previous section motivates the first research question (RQ):
  • 1.
    RQ.1: What kind of OSS have the STA and OpenSSF funded and thus considered as critical?
An answer to RQ.1 sheds important light on what kind of OSS is seen as critical by a government-backed agency and a not-for-profit OSS foundation in terms of cyber and software security. Indirectly, the question also aligns with existing work on how to identify critical infrastructures (Di Mauro et al.,, 2010; Katina and Hester,, 2013). In this regard, it can be also remarked that the OpenSSF, 2024e has a specific working group trying to solve the issue of how to identify critical open source software projects. For the present purposes, an answer to RQ.1 can be reflected also from a broader perspective; if some particular type of OSS stands out, it may be that the EU’s existing legal frameworks have overlooked some important aspects when operating with the criticality concept.
In this regard, a further EU law can be noted: Regulation (EU) 2024/2847, which is commonly known as the Cyber Resilience Act (CRA). Unlike in the NIS2 directive, criticality in this regulation is not approached through sectors but through products (Eckhardt and Kotovskaia,, 2023). It also specifies some new cyber security requirements for commercially supported OSS projects, including with respect to vulnerability disclosure and associated coordination (Ruohonen and Timmers,, 2024). Open source foundations, which often fund open source projects, are also mandated to develop policies for handling vulnerabilities and fostering secure software development. In addition, commercial vendors who integrate OSS into their products are mandated to notify the associated OSS projects about vulnerabilities they have discovered or been made aware of. Meeting the CRA’s essential cyber security requirements, which have been recently analyzed (Ruohonen et al.,, 2025), in the commercial software industry necessitates dealing also with OSS because many—if not most—commercial software projects use at least some OSS components. For instance, among the essential requirements is an obligation to ship products without known vulnerabilities. This legal requirement applies also to all OSS components embedded to a product.
While keeping in mind potential unintended consequences (Ruohonen,, 2024; Ryan et al.,, 2023), these legal mandates can be seen to implicitly address the noted interlinking aspect of critical infrastructures. The EU has also established new funding schemes to support the implementation and adoption of the new cyber security laws in the public and private sectors (Ruohonen and Timmers,, 2025; Veigurs et al.,, 2024). However, these funding instruments seem to exclude OSS projects; hence, an answer to RQ.1 can also help to understand the point about regulators and political decision-makers perhaps having overlooked important aspects related to critical infrastructures and their development, maintenance, and operation.
Regarding grants and funding in general, the following research question is worth asking:
  • 1.
    RQ.2: What have been the monetary amounts granted and the durations of the projects funded by the STA and OpenSSF?
The money involved sheds light on the overall importance of the two funding bodies for the OSS world in general. Regarding the funding durations, it has been suggested to separate short term (less than a year), medium term (from a year to three years), and long term (more than three years) funding from each other (Osborne et al.,, 2024). Against this backdrop, an answer to RQ.2 can potentially reveal insights on the cyber security issues involved in the OSS domain. If there have been a lot of short term funding but also some particular long term projects in this domain, it might be tentatively concluded that also the security challenges vary according to a context.
The types of OSS projects funded and the money involved can reveal a lot but obviously not everything. Particularly the wording in RQ.1 about criticality needs sharpening. By following existing guidelines regarding funding objectives (Osborne et al.,, 2024), the point about criticality motivates the subsequent research question:
  • 1.
    RQ.3: What have been the rationales of the STA and OpenSSF for granting funding to OSS projects?
The question can be motivated by the extensive research on choosing OSS components for a software project, whether commercial or open source. This research area is closely related to the survivability and sustainability aspects noted in the introduction. The list of factors to consider is extensive already in terms of security alone. The examples include reported vulnerability counts for a component (Cox,, 2019), the centrality of a component in a dependency network and the trust placed upon its maintainers (Zimmermann et al.,, 2019), and even recommender systems using these and other factors for recommendations (Nguyen et al.,, 2018). By and large, however, the existing research has operated with quantified metrics that can be easily extracted from OSS projects. To this end, an answer to RQ.3 addresses a notable limitation in existing research because it may well be that a rationale for funding has only a little to do with the conventional, mechanical metrics. This point entwines with previous observations according to which developers mostly concentrate on functionality and not security when choosing OSS components and dependencies (Pashchenko et al.,, 2020). Here too the incentives noted in the introduction are visible. In any case, an answer to RQ.3 can also be beneficial to OSS projects and their developers in formulating competitive funding proposals. In this regard, RQ.3 further aligns with existing research on guidelines for writing grant applications in the domain of software projects (Strasser et al.,, 2022). In addition to the types of software and the rationales for funding decisions, it is relevant to know what kind of cyber security the funded projects have addressed:
  • 1.
    RQ.4: What kind of cyber security the OSS projects funded by the STA and OpenSSF have addressed?
An answer to RQ.4 sheds further light on how the funding bodies have perceived criticality in terms of cyber security. In addition, the answer helps to better understand what kind of cyber security bottlenecks there are in the OSS domain according to the two funding bodies. However, it may well be that the real or perceived bottlenecks differ between the funding bodies. This point can be extended and presented as the final research question:
  • 1.
    RQ.5: Do the answers to RQ.1, RQ.2, RQ.3, and RQ.4 differ between the STA and OpenSSF?
An answer to the final RQ.5 serves two functions. Although the paper follows a qualitative methodology, the first function is an overall validity check. The second function is related to the strategic nature of the two funding bodies. If there are notable differences, it would seem reasonable to conclude that also the strategies of the funding bodies may differ. In fact, it could be even argued that the strategies should differ because overlapping funding could be interpreted to waste resources and highlight poor strategic foresight. On these notes, the materials and methods can be subsequently elaborated.
4 Materials and Methods
The data is based on all funding cases that were available from the OpenSSF, 2024a’s () and STA, 2024i’s () websites during the start of the data collection on 27 November, 2024. The data includes all funding cases listed on the websites, including those without longer funding descriptions. In total, $n=24$ and $n=44$ funding cases are covered from the OpenSSF and STA, respectively.
The paper’s qualitative methodology is based on thematic analysis. Somewhat like topic modeling in quantitative research, this classical qualitative method seeks to find latent themes present in a corpus. An inductive logic was used for the identification of themes and their refinement. By following guidelines (Terry et al.,, 2017; Vaismoradi et al.,, 2013), the thematic analysis was carried out in five steps. The first step was about familiarizing with the qualitative data, meaning that the funding cases were read and reread, and initial notes were taken. Based on this familiarizing, the second step involved a generation of initial codes for the potential themes. Then, the third and fourth steps were about searching and identifying themes based on the codes generated, and then collating overlapping themes due to a goal of parsimonious answers to the five research questions specified. The fifth and final step involved defining and naming the final themes. Although the result are summarized with concise, code-like themes together with their occurring frequencies, the common pitfalls of mere categorization and plain quantification (Saunders et al.,, 2023) are addressed by illustrating the themes with further elaborations and contemplations relevant to further research.
Regarding RQ.3 and RQ.4, three most important overall themes were identified for both research questions, provided that sufficient information was available. While these are the three most important themes in general, there is no order of importance among the three themes themselves. This choice can be argued to be necessary because many of the funding cases have addressed multiple distinct topical areas akin to work packages in scientific grant applications.
It can be noted that thematic analysis is sometimes triangulated with other qualitative methodologies, including what is known as grounded theory. However, a step toward deductive reasoning with prior theoretical premises was omitted on the grounds that the existing OSS funding research is too scant for developing sound theoretical building blocks. Though, regarding grounded theory, which essentially assumes that a theory can be found from data—it is grounded there, the final Section 6 discusses also theoretical implications of the qualitative results reached. Grounded theory provides also means to further elaborate the coding process behind the thematic analysis: the initial identification of themes involved open coding, which resulted in many themes, which were then refined and reduced by the means of axial coding, and selective coding was finally used for collating and sharpening the themes from the axial coding (Stol et al.,, 2016; Williams and Moser,, 2019). What separates the coding done from grounded theory is that no explicit theorization was done alongside the analysis and the themes were not taken as latent (cf. Terry et al.,, 2017). In addition to the procedural coding logic, it is important to emphasize that all the paper’s authors were involved in the analysis.
It is also worth emphasizing that the coding was rather straightforward due to the relatively short descriptions provided for the funding decisions. In general, the STA tends to provide a little more information than the OpenSSF, but even that information typically only amounts to a few short paragraphs. Both funding bodies describe their decisions in unstructured formats. In addition to the slight variance between the funding bodies, there is a variance between individual funding decisions; some are described shortly, while others are elaborated in detail. This variance justifies also the qualitative approach because a quantitative approach might be biased due to the variance. The simplicity of the final, collated themes eased the coding and its robustness. For instance, it is rather trivial to determine whether a funded project is about operating systems (e.g., OpenSSF, 2024b, ), logging libraries (STA, 2024d, ), desktop environments (STA, 2024c, ), or programming language ecosystems (e.g., OpenSSF, 2024c, ; STA, 2024b, ). As can be deduced from Fig. 1, most of the funding decisions and the themes for these do not require complicated reasoning.
Figure 1: An Example of a Funding Decision (STA, 2024d, )
Thus, to further elaborate the collaborative side of the thematic analysis, the initial analysis was carried out by the paper’s first author. Then, the paper’s other collaborating authors independently reviewed, case by case, the results from the initial analysis. They were given a chance to disagree and propose corrections, including the introduction of new themes. With respect to disagreements and diverging themes, the first author made the final decision about a theme by evaluating the reviews of the collaborating authors. Although inter-rater reliability measures were omitted, this collaborative process follows the recent suggestion to work sequentially in qualitative coding (Díaz et al.,, 2023). The collaborative approach is also known as researcher triangulation in qualitative research (Farquhar et al.,, 2020). Regarding inter-rater reliability measures, the omission of these can be justified on the grounds that only the final themes are relevant and thus reported, the prior knowledge of the authors varied regarding the subject matter, and the funding decisions were rather straightforward to interpret and categorize (cf. McDonald et al., 2019; and more generally Terry et al., 2017 who argue against the measures altogether in qualitative research). Even though independent and blinded coding was not done, the collaborative approach can be argued to still improve the rigor and trustworthiness of the qualitative results. Though, with the earlier Fig. 1 in mind, only a few decisions prompted some initial disagreements.
5 Results
The dissemination of the qualitative results can be started by noting that ten out of the $68$ funding cases did not address cyber security. A good example of non-security funding would be a grant for the GNOME desktop environment (STA, 2024c, ). These cases are thus excluded from the answers given to RQ.4. Then, regarding RQ.1, a concise thematic summary is presented in Table 1 about the types of OSS funded by the two funding bodies.
As can be seen, open source software supply chains, including package managers for installing, updating, and maintaining dependencies, have been the most frequently funded target. The result is not surprising because the security issues with OSS supply chains are well-recognized (Hamer et al.,, 2025; Ohm et al.,, 2020; Ruohonen and Saddiqa,, 2025; Zimmermann et al.,, 2019). These issues affect large user and deployment bases, including in the commercial software industry due to the widespread use of OSS components and dependencies. Supply chain security and management are also more generally seen as being among the foremost difficult global issues in the increasingly globalized and digitalized world (Wu,, 2024). Thus, it is no wonder that the STA and OpenSSF too have recognized their criticality.
Table 1: Types of OSS Funded
Theme Frequency
Supply chains and package managers 25
Network libraries 12
Programming languages 8
Operating systems and their components 6
Cryptographic libraries (PGP)1 4
Scientific computing 3
Multimedia libraries 3
Software testing frameworks 2
Logging libraries 2
Web servers 1
Web frameworks 1
Desktop environments 1
$\sum$ 68
  • 1.
    1 Pretty good privacy (PGPs).
The security of OSS supply chains also aligns with the funding granted to programming language projects because practically all major programming languages nowadays have their own ecosystems for distributing packages. A further point is that the funding for improving OSS supply chain security aligns with the noted CRA regulation. The reason is that particularly the OpenSSF has funded other open source software foundations that typically govern the ecosystems and their supply chains. In the CRA these are known as “stewards”. Although only a lighter compliance burden is imposed upon them by the CRA, they are still mandated to improve particularly vulnerability disclosure and associated vulnerability coordination through establishing policies, promoting sharing of vulnerability information, and collaborating with European market surveillance authorities. Thus, it can be argued that the funding from the STA and OpenSSF have patched a limitation in a sense that the EU’s own funding instruments have largely excluded the so-called stewards. Further research would be needed to deduce whether the CRA was strategically framed during the political negotiations with respect to this funding issue. In the meanwhile, it can be concluded that the cyber security funding for OSS projects aligns also with new regulatory demands.
Network libraries and protocol implementations have also been frequently funded particularly by the STA. As could be expected, DNS and BGP implementations and their projects are among the funding recipients. Also other networking libraries and implementations crucial for the whole Internet are among the recipients, including those implementing the transport layer security (TLS) and secure shell (SSH) cryptographic protocols (OpenSSF, 2024d, ; STA, 2024f, ). It is worth mentioning also funding for a project implementing end-to-end encryption protocols as well as four separate funding decisions for projects implementing the pretty good privacy (PGP) cryptography protocol. Of the newer networking protocols related to security, WireGuard has also received funding.
Finally, regarding the concept of criticality, it is understandable that also operating systems and their low-level software components have received funding. An example would be FreeBSD (OpenSSF, 2024b, ), but the OpenSSF has also granted funding to the Linux kernel, which is closely associated with the OpenSSF’s parent organization, the Linux Foundation. Although the data is too limited for answering to the associated why-question, this observation is enough to remark that the OSS funding bodies—like many but not necessarily all funding bodies and financing instruments in general, may not be entirely immune to potential biases and conflicts of interests. A further point is that criticality as a concept seems to align with the STA, 2024i’s () notion of foundational technologies without which the Internet may function poorly and the creation of new software may become difficult. The frequent funding of OSS supply chains further aligns with the OpenSSF, 2024e’s () emphasis of dependencies for defining criticality. However, there is also some room for criticism regarding the concept of criticality and its relation to the funding decisions. For instance, STA, 2024g has granted $450$ thousand euros for a project seeking to improve documentation about a JavaScript library. While documentation is important, it remains debatable how critical the library is with respect to the foundational technologies emphasized by the STA.
Turning to RQ.2, the monetary amounts granted are visualized in Fig. 2. (It can be noted that at the time of writing the exchange rate between the EU’s euro and the United States dollar was almost even, and thus no conversion was needed.) The average is a little below four thousand euros, but there are a few outlying projects that have received over a million euros. Without prior work, it is rather difficult to contemplate how large or small these amounts are. Nevertheless, consider an imaginary example of a professional software engineer earning roughly around five thousand euros a month; then, the average funding amount would already account for about six full-time software engineers working with a project for a year. This example allows to argue that the funding amounts are decent enough in the OSS domain. Although the OpenSSF does not provide data on the funding durations, most (84%) of the STA’s funding decisions have been for two years. This observation further reinforces the point about the interpretation of the money involved.
Figure 2: Funding Amounts
Table 2: Types of Cyber Security Funded
Theme Frequency
STA OpenSSF Both
Supply chain security1 16 17 33
Cryptography 9 0 9
Network security 8 0 8
Security testing 4 0 4
Memory safety 4 2 6
SBOMs 3 1 4
Maintenance 3 2 5
Audits and auditing 3 11 14
Security education 1 2 3
APIs2 1 0 1
$\sum$ 52 35 87
  • 1.
    1 Excluding software bills of materials (SBOMs).
  • 2.
    2 Application programming interfaces (APIs).
The types of cyber security funded (RQ.4) summarized in Table 2 expectedly align closely with the earlier thematic summary in Table 1. In other words, supply chain security has been the most typical thematic cyber security area funded. In this regard, it is important to further note that some of the OSS projects funded have also addressed the design and implementation of so-called software bills of materials (SBOMs). These materials for keeping track of dependencies and other components are noteworthy for two reasons, the first being that SBOMs are also under active research in software engineering (Bi et al.,, 2024; Xia et al.,, 2023, among others). The second reason is that SBOMs too are present in the EU’s CRA, which mandates commercial software vendors to implement SBOMs for their products and deliver these to market surveillance authorities upon request (Ruohonen and Timmers,, 2024). Although OSS projects and their so-called stewards are excluded from this regulatory requirement, the integration of OSS components into commercial software products might partially explain also the funding decisions for SBOMs by the STA and OpenSSF. Another partial explanation might be that also the funding bodies perceive SBOMs as important for improving supply chain security and software security of OSS in general.
Then, regarding RQ.5, already the summary in Table 2 is enough to conclude that there is a divergence between the funding bodies also with respect to the types of cyber security funded. In particular, the STA has funded also many projects addressing encryption and cryptography, network security, and security testing, whereas the OpenSSF has been particularly active in funding of security audits, including those done by third parties. Audits and auditing are important to underline because these have involved also commercial cyber security companies doing the auditing. Thus, the money involved may sometimes go to the private sector even though the goal of funding remains the same. Furthermore, audits and auditing align with the goals of the CRA regulation.
There are also some projects that have received funding from both OSS funding bodies. The examples include FreeBSD and the Python Software Foundation, which is maintaining the popular and large repository for distributing software packages written in Python. The data is again too limited for far-reaching speculations, but, nevertheless, an argument can be raised that the two funding bodies might benefit from joint coordination. This point can be strengthened by noting that the data contains also some cases indicating joint ventures between different funds (OpenSSF, 2024d, ). Against this backdrop, it might be even reasonable to argue that the whole OSS funding landscape might benefit from a high-level coordination body through which all funds could collaborate.
Table 3: Funding Rationales
Theme Frequency
STA OpenSSF Both
Cyber security 19 23 42
Sustainability 18 1 19
New features 14 3 17
Testing and bug fixing 13 11 24
Documentation and education 9 2 11
Performance 7 0 7
Interoperability 7 0 7
Rewrites 4 2 6
Infrastructure 4 0 4
Usability 3 0 3
Reliability 3 0 3
Auxiliary tools 3 0 3
Software engineering processes 1 13 14
Compliance and assurance 1 0 1
Accessibility 1 0 1
$\sum$ 52 35 87
With respect to RQ.3, the funding rationale themes are summarized in Table 3. To put aside cyber security itself, which has sometimes been justified with explicit mentions of critical infrastructure (STA, 2024b, ), also the sustainability theme is strongly present as a rationale. Sustainability as a theme also aligns with the rather frequent rationales of software testing and plain old bug fixing. However, it is worthwhile to mention that some projects have approached the sustainability conundrum through identifying and then deprecating components and dependencies that have reached their end-of-life software life cycle stages or became technologically risky (STA, 2024e, ; STA, 2024f, ). In a similar vein, the earlier point about technological debt is visible in the few projects that have approached sustainability problems through wholesale rewrites and re-implementations. Such solutions are often interpreted as a sign of technical debt of the most severe kind (Verdecchia et al.,, 2021). However, the rewrites and re-implementations have not been only about technological debt and sustainability in general.
The reason is that some of these have involved also transformations to other programming languages; hence, the rewrites and re-implementations also align with the memory safety theme in the earlier Table 2. The Rust programming language is the prime example in this regard. The language and its ecosystem have also received funding from both bodies. From the STA, 2024h the Rust project received funding for implementing SBOMs, while from the OpenSSF, 2024c it received funding for a myriad of security themes, including hiring a dedicated security engineer, audits, documentation, advocacy, and education, as well as new security tools and an establishment of a yet another grant program for mentoring and education. The Rust project’s own grant program yet again highlights the potential benefits from better coordination. The OpenSSF, 2024c’s () funding case also indicates that the funding applications and their evaluations might benefit from a more structured approach. For instance, like in large-scale scientific grant applications, it could be contemplated whether structuring the themes and rationales into work packages would bring more structure and coherence. A further point is that the education, outreach, and documentation involved, and the accompanying overall theme in Table 3, again frames the OSS funding topic also toward academia. In this regard, it could be argued that the security themes funded by the STA and OpenSSF are also something that should be taught in higher education. The Rust case also demonstrates that there might be a need in the nearby future to update curricula with respect to programming languages.
Finally, it is important to emphasize that the funding rationales have involved also implementing new features explicitly or implicitly related to software security and cyber security in general. In addition, also many other general software engineering themes are present, including performance and scalability, interoperability, reliability, usability, and software engineering processes and practices. These rationales allow to end the qualitative exposition by arguing that the tragedy of the (digital) commons alone can neither explain nor elaborate the whole topic of OSS funding. Also, many other explanations are presumably present. The PPP aspect noted is among these explanations; the private sector also greatly benefits from OSS. Thus, further research is needed to better understand the business aspects behind OSS funding. Concepts such as so-called coopetition (Duc et al.,, 2017) might provide a theoretical path to follow in this regard.
6 Discussion
6.1 Conclusion and Contributions
The paper presented a qualitative overview of cyber security funding granted to various OSS projects through two important funding bodies. Both bodies fund OSS projects that are seen as being closely related to the world’s critical infrastructures. Among these software-defined infrastructures are OSS supply chains, core computer networking libraries, servers, operating systems and their low-level software components, and programming languages, among other things. The thematic cyber security areas funded are closely related; supply chain security, cryptography, network security, memory safety, audits, and security education, among other things. Both funding bodies connect improvements at these areas to the sustainability problems faced by many OSS projects.
However, cyber security and sustainability alone cannot fully explain the rationales for funding; therefore, neither is the tragedy of the (digital) commons sufficient as an overall theoretical explanation. In addition to raising this important argument, the paper made a notable contribution by linking together the research branches addressing the sustainability of OSS projects and critical infrastructure. Regarding the latter research branch, software’s role has often been downplayed to some extent, and the role of OSS has been almost nonexistent in existing research. A further contribution was made by elaborating how the OSS cyber security funding theme examined is related to regulations. Particularly the EU’s new CRA regulation is relevant also for open source software projects. Against this backdrop, the paper further contributed to recent efforts to better understand the EU’s new funding instruments (Ruohonen and Timmers,, 2025). It seems that the two funding bodies examined have patched either intentional or unintentional omissions in more general public sector funding in Europe and likely also elsewhere. In a similar vein, also threat intelligence reports about supply chain security have missed the sustainability of OSS projects as a mitigative factor (Hamer et al.,, 2025). Finally, the paper contributed to the still nascent theorization and framing of funding for open source software projects.
The point about theorization and framing deserves three additional brief remarks. The first remark is that the involvement of the private sector aligns the OSS funding topic toward so-called public-private partnerships (PPPs), which have been common in the cyber security domain in general (Ruohonen,, 2020). Insofar as OSS is interpreted to be on the public side, which would seem a reasonable interpretation with respect to the concept of public goods underneath the tragedy of the commons, PPPs are particularly visible in the OpenSSF case because the associated fund was established by donations from large technology companies, among them Amazon, Google, and Microsoft (OpenSSF,, 2023). The second remark follows: PPPs are typically also required for funding of large-scale applied scientific projects, including those funded by the EU’s Horizon Europe program. These two remarks logically lead to the final remark: the OSS funding topic is also about the so-called technology transfer model, which has been endorsed and debated in software engineering research for a long time (Mikkonen et al.,, 2018; Wohlin et al.,, 2024, among many others). Regarding the transfers involved, however, scientific research and academia are missing. Given the Germany’s involvement behind the STA, the transfers are primarily between the public sector, the public (OSS) domain, and the private sector.
Therefore, it could be contemplated whether adding academic software engineering research to the mixture would improve the security of OSS even further, perhaps opening also new collaborative avenues and further funding possibilities. This contemplation is reinforced by the educational efforts done by some OSS projects with the help of funding from the two bodies. There are also some indications in the data that some projects have already justified their funding applications with references to scientific research projects and the EU’s own software development projects (STA, 2024e, ). Thus, the earlier points about coordination can be extended toward a larger governance space, which, in Europe, involves also the cyber security regulations mentioned.
6.2 Further Work
The concept of criticality is the keyword behind the funding decisions. However, as was discussed, it is not entirely clear how the two funding bodies have used the concept in decision-making about funding applications. For instance, the corresponding working group established by OpenSSF, 2024e has emphasized dependencies, popularity, and related concepts in determining a project’s criticality. However, dependencies and popularity may align poorly with the STA’s notions about open digital infrastructures and foundational technologies. Consider DNS and BGP implementations as an example; they are clearly critical for the Internet’s functioning, but they may not be as popular as some other OSS projects with large user bases. Nor do they garner a large amount of dependencies. Thus, further research is required about the criticality concept and its practical use in funding decisions.
In addition, an important point regarding further research is the evaluation of not only funding applications but the successes or failures of the projects funded. In this regard, OpenSSF, (2023) has noted that their evaluation has been “largely anecdotal” and that further goals include “highlighting the success factors to future grant recipients and using them in our grant review process”. The STA, 2024a’s () evaluation criteria for funding applications are a little more systematic; prevalence, relevance, sustainability, public interests, activities planned and their feasibility, and expertise of the applicants are evaluated. However, the Agency’s online materials do not say anything about whether funded OSS projects are evaluated post hoc after their completions. The evaluation criteria and related points could be investigated in further research by means of interviews, focus groups, or even surveys, given the apparent proliferation of different funds for OSS projects.
Different evaluation criteria have been considered and compared also in existing research (Osborne et al.,, 2024), but it remains unclear how applicable they are in practice. Science funding provides a good example. On one hand, learning from the funding of scientific projects might help with the goal of improving evaluations. Provided that the point about potential biases and conflicts of interests is true even to some extent, the learning might offer means to improve objectivity in reviewing because objectivity has traditionally been highly valued in science. For instance, it is unclear whether the STA and OpenSSF have used external reviewers, and whether such reviewers would be seen as preferable. On the other hand, it could be also argued that science funding might be a poor reference because it is often rather bureaucratic.
That is, lengthy and rigorous applications as well as extensive reporting and accounting might well conflict the values and practices endorsed by OSS communities. Analogous conflicts have been reported also in existing research with respect to some OSS communities perceiving sponsors’ and stakeholders’ interests and demands as being in a conflict with community values, norms, and project goals (Osborne,, 2024). The STA and OpenSSF cases are also good examples in a sense that the overall sustainability theme aligns only poorly with the novelty theme usually strongly present in science funding. Furthermore, as has been debated for a long time (Herley and Van Oorschot,, 2017), the other main theme, cyber and software security, is difficult to evaluate by empirical means.
Regardless whether science funding should or should not be used for learning, it provides many themes for formulating relevant research questions. For instance, a classical result has been that winning grants increases the probability of winning more grants in the future (Bol et al.,, 2018; Mocanu et al.,, 2024). It would be interesting to know whether a similar effect applies in the OSS funding context. The STA, 2024a’s () emphasis on expertise hints that a similar effect might be present at least to some extent. The question also correlates with the evaluation theme; it remains generally unclear what makes a competitive funding application in this context. Due to the varying domains of OSS and the numerous distinct funds, it also remains unclear whether the guidelines presented in existing research (Strasser et al.,, 2022) generalize. A further question that needs addressing is related to the awareness among OSS communities about the new funding opportunities available. It may be that the STA’s and OpenSSF’s funding decisions have been biased not because of evaluations but because not all critical OSS projects have been aware of the funds. Finally, as was discussed and argued, it should be evaluated whether and how academic research might join to future funding applications. This final point can be also used to argue that the technology transfer model may need some altering, updating, and theorizing.